Find the general solution of the following differential equations using method of undetermined coefficients: (iii) y''+ 4y' = sin5x+x−1,


1
Expert's answer
2022-03-01T13:46:33-0500

"y''+4y'=\sin 5x+x-1\\
y=y_1+y_2\\
1) y''+4y'=0\\
\lambda^2+4\lambda=0\\
\lambda_1=0, \lambda_2=-4\\
y_1=c_1+c_2 e^{-4x}\\
2) y_2=a \cos 5x+ b \sin5x +x(kx+p)\\
y'_2 =-5a \sin5x +5b \cos5x +2kx+p\\
y_2''=-25a \cos5x-25 b\sin5x+2k\\
-25a \cos5x-25 b\sin5x+2k -20a \sin5x +\\
+20b \cos5x +8kx+4p=\sin 5x+x-1\\
\cos5x: -25a+20b=0\\
\sin5x: -25b-20a=1\\
x: 8k=1\\
x^0: 2k+4p=-1\\
a=-\frac{4}{205}\\
b=-\frac{1}{41}\\
k=\frac{1}{8}\\
p=-\frac{5}{16}\\
y_2=-\frac{4}{205} \cos 5x-\frac{1}{41} \sin5x +x(\frac{1}{8}x-\frac{5}{16})\\
y=c_1+c_2 e^{-4x}-\\
-\frac{4}{205} \cos 5x-\frac{1}{41} \sin5x +x(\frac{1}{8}x-\frac{5}{16})"